A vehicle tire has a radius of 0.27 meters. If the vehicle accelerates in a straight line from rest at 2.85 m/s/s, calculate for the angular acceleration of the front passenger tire.
1
Expert's answer
2020-11-23T05:36:20-0500
"\alpha=\frac{a}{r}\\\alpha=\frac{2.85}{0.27}=10.6\frac{rad}{s^2}"